OpenStudy (toxicsugar22):

Find the first three terms of the product in acending powers. (1-2x^2-x^4)(4+x^2-2x^4)

OpenStudy (toxicsugar22):

do u undestand

OpenStudy (sasogeek):

pretty much, just figuring out a couple of stuff

OpenStudy (sasogeek):

you know what, i'm not going to waste loads of time with formulas. I'm not an expert but i'll tell you the obvious. find the the first 3 terms(a,b,c) of the product (ef) in acending powers.

OpenStudy (sasogeek):

simple mutiplication.... \[= 1(4+x ^{2}-2x ^{4}) - 2x ^{2}(4+x ^{2}-2x ^{4})-x ^{4}(4+x ^{2}-2x ^{4})\] \[=4+x ^{2}-2x ^{4}\] ^^^^^^^^^^^ First 3 terms

OpenStudy (toxicsugar22):

ok thanks

OpenStudy (sasogeek):

I'm not too sure but its just what i think out of the obvious